OpenStudy (anonymous):

Please help. Check my answer (commented below). Kim wants to have equal amounts of room on all four sides of her dining room table. A 5ft by 8ft table is placed in the center of a rectangular room that has an area of 154 square feet. What is the width of the clear space Kim will have on all sides of her table?

OpenStudy (anonymous):

Is this right?? x=width Length of room=8+2x Width=5+2x (8+2x)(5+2x)=154 40+10x+16x+4x^2=154 4x^2+26x-114=0 2x^2+13x-57=0 2x^2-6x+19x-57=0 2x(x-3)+19(x-3)=0 (2x+19)(x-3)=0 x=-19/2 x=3

OpenStudy (saifoo.khan):

Seems right.

OpenStudy (anonymous):

Thanks!

OpenStudy (saifoo.khan):

You can check your answer by inserting the value "3" in your expression.

OpenStudy (saifoo.khan):

(8+2x)(5+2x)=154 [8+2(3)] * [ 5+2(3)] Solve and see if you get 154.

OpenStudy (anonymous):

Got it thanks :)